OpenStudy (anonymous):

A(n) 76.6 kg astronaut becomes separated from the shuttle, while on a space walk. She finds herself 54.9 m away from the shuttle and moving with zero speed relative to the shuttle. She has a(n) 0.545 kg camera in her hand and decides to get back to the shuttle by throwing the camera at a speed of 12 m/s in the direction away from the shuttle. How long will it take for her to reach the shuttle? Answer in minutes. Answer in units of min

OpenStudy (anonymous):

momentum is the same' 76.6*v=0.545*12 find v \[Then time=\frac{ distance }{v }\]